The Empire Hotel is a full-service hotel in a large city. Empire is organized into three departments that are treated as investment centers. Budget information for the coming year for these three departments is shown as follows. The managers of each of the departments are evaluated and bonuses are awarded each year based on ROI.

Empire Hotel
Hotel Rooms Restaurants Health Spa
Average investment $ 8,247,000 $ 4,802,000 $ 848,000
Sales revenue $ 10,000,000 $ 2,000,000 $ 600,000
Operating expenses 8,682,000 1,390,000 486,000
Operating earnings $ 1,318,000 $

610,000

$ 114,000

Required:

a. Compute the ROI for each department. Use the DuPont method to analyze the return on sales and capital turnover.

Assume the Health Spa is considering installing new exercise equipment. Upon investigating, the manager of the division finds that the equipment would cost $40,000 and that operating earnings would increase by $8,000 per year as a result of the new equipment.

b-1. What would be the ROI of investment in the new exercise equipment and Health Spa?

b-2. Would the manager of the Health Spa be motivated to undertake such an investment?

c-1. Compute the residual income for each department if the minimum required return for the Empire Hotel is 17 percent.

c-2. What would be the impact of the investment on the Health Spa's residual income?
Answer #1

Part a

Departments

ROI

Return on sales

Capital turnover

Hotel Rooms

15.98%

13.18%

1.21

Restaurants

12.70%

30.50%

0.42

Health Club

13.44%

19.00%

0.71

Empire Hotel

14.69%

16.21%

0.91

ROI = operating earnings / average investment

Hotel Rooms = 1318000/8247000 = 15.98%

Restaurants = 610000/4802000 = 12.70%

Health Club = 114000/848000 = 13.44%

Empire Hotel = 2042000/13897000 = 14.69%

Return on sales = operating earnings / sales revenue

Hotel Rooms = 1318000/10000000 = 13.18%

Restaurants = 610000/2000000 = 30.50%

Health Club = 114000/600000 = 19.00%

Empire Hotel = 2042000/13897000 = 16.21%

Capital turnover = sales revenue / average investment

Hotel Rooms = 10000000/8247000 = 1.21

Restaurants = 2000000/4802000 = 0.42

Health Club = 600000/848000 = 0.71

Empire Hotel = 2042000/13897000 = 0.91

Part b-1

ROI of investment in the new exercise equipment and Health Spa = (114000+8000)/(848000+40000) = 13.74%

Part b-2

Yes, the manager of the Health Spa would be motivated to undertake such an investment as there is increase in ROI

Part c-1

Departments

Residual income

Hotel Rooms

-83990

Restaurants

-206340

Health Club

-30160

Residual income = operating earnings – return on investment

Hotel Rooms = 1318000-(8247000*0.17) = -83990

Restaurants = 610000-(4802000*0.17) = -206340

Health Club = 114000-(848000*0.17) = -30160

Part C-2

There will be positive impact of the investment on the Health Spa's residual income and the residual income would increase by $1200 (8000-(40000*17%))